OpenStudy (anonymous):

Data set X and data set Y both have the same interquartile range. If the first quartile of data set X is less than the first quartile of data set Y, then the third quartile of data set X is greater than the third quartile of data set Y. True or False?

OpenStudy (mathmate):

Consider the data sets: X={1,1,3,4,5,5,5}, IQR=5-1=4 Y={1,2,3,4,5,6,7}, IQR=6-2=4